Use the fact that the speed of light in a vacuum is about 3.00  108 m/s to determine how many kilometers a pulse from a laser beam travels in exactly 2 hour(s)


1
Expert's answer
2021-09-12T19:23:33-0400

"2hrs\times\frac{3600sec}{1hr}\times\frac{3.00\times10^8m}{1sec}\times\frac{1km}{1000m}=2.16\times10^9km"


Answer: "2.16\times10^9km"